OpenStudy (anonymous):

Solve for g: s=1/2gt^2

OpenStudy (tkhunny):

Have you considered multiplying by 2 and dividing by t^2? Change your attitude and get a different screen name.

OpenStudy (anonymous):

@tkhunny @Mokeira I tried multiple usernames and all of them were taken and this is one happened to be available so I used it. You don't have to be rude about it.

OpenStudy (asevilla5):

Multiply through by 2 to rid the fraction. 2s = gt^2 Divide both sides by t^2 ========> g = 2s / t^2

OpenStudy (anonymous):

@texaschic101 thanks so much for helping

OpenStudy (texaschic101):

or just divide both sides by 1/2 t^2 and you get : s = 1/2gt^2 -- divide by 1/2t^2 s/(1/2)t^2 = g

OpenStudy (inkyvoyd):

It's funny when the moderator is less polite than the user hahaha

OpenStudy (texaschic101):

so true....maybe it has just been a bad day for tkhunny

OpenStudy (inkyvoyd):

no, tkhunny is just like that

OpenStudy (asevilla5):

some people just don't know how be polite in the first place

OpenStudy (inkyvoyd):

besides, a moderator's job is to moderate, and by that kind of notion s/he's shifted the mood of the chat to some bluntness lol

OpenStudy (inkyvoyd):

I'm not saying I'm any better, it's just that I find it amusing that we're all.. human? idk

OpenStudy (inkyvoyd):

even the tkbanhamma

OpenStudy (anonymous):

Hi @texaschic101 I have two questions similar to the one you just answered. Do you think you could possibly help me again? Anyone who helps would be appreciated actually.

OpenStudy (inkyvoyd):

@xxoidontcarexxo just post a new question and tag the user.

OpenStudy (anonymous):

Okay thanks so much @inkyvoyd

OpenStudy (tkhunny):

My purpose is to help each student succeed. If your screen name suggests you cannot succeed, I will suggest that you change it. I am truly sorry that my honesty and directness are so often mistaken for rudeness. You can take it for what it is, instead, rather than construe rudeness. Anyone with a user name including "I Don't Care" or "I suck at math" or "I hate this" or quite a few other variations should change user names. Simple as that. No rudeness at all.